OpenStudy (anonymous):

simplify 3^ 1/2 * 3^ 1/2

OpenStudy (anonymous):

Well 3 1/2 =7/2 So 7/2 *7/2 = 49/4 49/4 is answer

OpenStudy (anonymous):

the question looks like this though. It makes it look like 3 to the power of 1/2 times 3 to the power of 1/2?

OpenStudy (anonymous):

3 1/2 * 3 1/2 = (3 1/2)^2

OpenStudy (jadeishere):

I've never seen that before. I got 2.25... I think i'm wrong

OpenStudy (anonymous):

yeah thats why Im so confused.

OpenStudy (jadeishere):

It's 12.25. 3 1/2 = 3.5 so 3.5 * 3.5 = 12.25 so 12 and 1/4

OpenStudy (anonymous):

@caters I still dont think thats right...im not just multiplying a regular fraction

OpenStudy (anonymous):

yeah but it's not 3 1/2 it's 3^1/2

OpenStudy (anonymous):

but the way you wrote it it looks like you are multiplying 3 + 1/2 by itself or squaring it.

OpenStudy (anonymous):

so 3*2=6+1=7 then divide that by 2 which is 3.5 so the answer is 7?

OpenStudy (anonymous):

thats still just adding a regular fraction...

OpenStudy (anonymous):

so you mean to say you are multiplying \[\sqrt{3}*\sqrt{3} = 3\]

OpenStudy (anonymous):

The answer is 3. I figured it out. Thanks for the help though everyone!